# Twisted character formula for toral supercuspidal representations

### v1

@@ 6.3. Descent of the Heisenberg quotient
Recall that $m_{\alpha}$ is the smallest positive integer such that $\theta^{m_{\alpha}}(\alpha)\in\Sigma\alpha$. We fix an element $\sigma_{\alpha}\in\Gamma_{\pm\alpha_{\mathrm{res}}}/\Gamma_{\alpha}$ as follows:
@@
    When $\alpha$ is asymmetric, we choose $\sigma_{\alpha}$ such that $\theta^{m_{\alpha}}(\alpha)=\pm\sigma_{\alpha}(\alpha)$. Note that, as $\alpha$ is asymmetric, exactly one of $\theta^{m_{\alpha}}(\alpha)$ or $\theta^{m_{\alpha}}(\alpha)$ belongs to $\Gamma\alpha$. In other words, the sign $\pm$” in $\theta^{m_{\alpha}}(\alpha)=\pm\sigma_{\alpha}(\alpha)$ is determined by $\alpha$.
@@
    When $\alpha$ is symmetric, we have $\theta^{m_{\alpha}}(\alpha)\in\Sigma\alpha=\Gamma\alpha$. Noting this, we choose $\sigma_{\alpha}$ such that $\theta^{m_{\alpha}}(\alpha)=\sigma_{\alpha}(\alpha)$.
@@
Let us show (1); suppose that $\alpha$ is asymmetric. We take any element $\sigma\in\Gamma_{\pm\alpha_{\mathrm{res}}}$ and suppose that $\sigma(\alpha)=\pm\theta_{\mathbf{S}}^{i}(\alpha)$ for some $i$. Then, as $m_{\alpha}$ is the smallest positive integer satisfying $\theta_{\mathbf{S}}^{m_{\alpha}}(\alpha)\in\Sigma\alpha$, we have $i\equiv jm_{\alpha}\pmod{l_{\alpha}}$ for some $j\in\mathbb{Z}$. Hence we have $\sigma(\alpha)=\sigma_{\alpha}^{j}(\alpha)$. This implies that $\sigma^{-1}\sigma_{\alpha}^{j}$ belongs to $\Gamma_{\alpha}$. In other words, $\sigma_{\alpha}$ generates $\Gamma_{\alpha_{\mathrm{res}}}/\Gamma_{\alpha}$.
@@
$V_{\eta}\cong\bigoplus_{\Sigma\alpha_{\mathrm{res}}\in\ddot{\Xi}_{\eta}}V_{\eta,\Sigma\alpha_{\mathrm{res}}},$
@@
$V_{\eta,\Sigma\alpha_{\mathrm{res}}}\cong\Bigl(\bigoplus_{\Sigma\alpha^{\prime}\in\Sigma\backslash(\Sigma\times\Theta)\alpha}V_{\Sigma\alpha^{\prime}}\Bigr)^{\eta}$
for any $\alpha_{\mathrm{res}}\in\Xi_{\eta}$. Let us put $V_{\Sigma\Theta(\alpha)}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\bigoplus_{\Sigma\alpha^{\prime}\in\Sigma\backslash(\Sigma\times\Theta)\alpha}V_{\Sigma\alpha^{\prime}}$. In particular, by letting $\Xi_{\mathrm{res}}$ be the set of restricted roots associated to $\Xi$, the set $\Xi_{\eta}$ can be thought of as the set of restricted roots $\alpha_{\mathrm{res}}\in\Xi_{\mathrm{res}}$ such that the $\eta]-$action has a nonzero fixed point in $V_{\Sigma\Theta(\alpha)}$.
@@ 6.4. Twisted characters of Weil representations: asymmetric roots
Let $\alpha\in\Xi_{\mathrm{asym}}$. We compute $\Theta_{\omega_{\Sigma\alpha}}([\eta]^{m_{\alpha}})$, which constitutes the right-hand side of Corollary 6.6. Recall that $V_{\Sigma\alpha}=V_{\Gamma\alpha}\oplus V_{-\Gamma\alpha}\cong V_{\alpha}\oplus V_{-\alpha}$, where $V_{\alpha}$ and $V_{-\alpha}$ are $1$-dimensional $k_{\alpha}$-vector spaces, which are identified with $k_{\alpha}$ by fixing nonzero elements $X_{\alpha}\in V_{\alpha}$ and $X_{-\alpha}\in V_{-\alpha}$ (see Section 6.1).
We carry out a case-by-case computation on the symmetry of $\alpha_{\mathrm{res}}$. Recall that we have exactly one of $\theta^{m_{\alpha}}(\alpha)\alpha\in\Gamma\alpha$ or $\theta^{m_{\alpha}}(\alpha)\alpha\in-\Gamma\alpha$; we have fixed a $\sigma_{\alpha}$ satisfying $\theta^{m_{\alpha}}(\alpha)=\sigma_{\alpha}(\alpha)$ or $\sigma_{\alpha}(\alpha)$ according to it. In the former case where $\theta^{m_{\alpha}}(\alpha)=\sigma_{\alpha}(\alpha)$, $\sigma_{\alpha}$ belongs to $\Gamma_{\alpha_{\mathrm{res}}}$, which implies that $\alpha_{\mathrm{res}}$ is asymmetric. On the other hand, in the latter case where $\theta^{m_{\alpha}}(\alpha)=-\sigma_{\alpha}(\alpha)$, the Galois group of $F_{\alpha}/F_{\alpha_{\mathrm{res}}}$ is a proper subgroup of that of $F_{\alpha}/F_{\pm\alpha_{\mathrm{res}}}$ generated by $\sigma_{\alpha}^{2}$, which implies that $\alpha_{\mathrm{res}}$ is symmetric. Thus, also recalling Lemma 6.8, we have the following possibilities:
@@ 6.4.1. The case where $\alpha_{\mathrm{res}}$ is asymmetric
The action of $\eta]^{m_{\alpha}}$ on $V_{\Sigma\alpha}$ preserves $V_{\Gamma\alpha}$ and $V_{-\Gamma\alpha}$. Since $\eta]^{m_{\alpha}}(X_{\alpha})$ belongs to $V_{\theta^{m_{\alpha}}(\alpha)}=V_{\sigma_{\alpha}(\alpha)}$, the induced action of $\eta]^{m_{\alpha}}$ on $V_{\alpha}$ is $\varsigma_{\alpha}$-linear:
${V_{\alpha}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${V_{\Gamma\alpha}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{[\eta]^{m_{\alpha}}}$${X_{\alpha}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\sum_{\sigma\in\Gamma/\Gamma_{\alpha}}\sigma(X_{\alpha})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${V_{\alpha}}$${V_{\Gamma\alpha}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${[\eta]^{m_{\alpha}}\circ\varsigma_{\alpha}(X_{\alpha})}$${\sum_{\sigma\in\Gamma/\Gamma_{\alpha}}[\eta]^{m_{\alpha}}\circ\sigma(X_{\alpha})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
@@
as an element of $\operatorname{Sp}(V_{\Sigma\alpha})\cong\operatorname{Sp}(k_{\alpha}\oplus k_{\alpha})$. Hence, by Corollary A.8, we get
$\Theta_{\omega_{\Sigma\alpha}}([\eta]^{m_{\alpha}})$ $=\operatorname{sgn}_{\mathbb{F}_{p}^{\times}}(\det(\eta_{\alpha}\circ\varsigma_{\alpha}\mid k_{\alpha}))\cdot|V_{\Sigma\alpha}^{\eta^{m_{\alpha}}}|^{\frac{1}{2}}$
$=\operatorname{sgn}_{\mathbb{F}_{p}^{\times}}(\det(\varsigma_{\alpha}\mid k_{\alpha}))\cdot\operatorname{sgn}_{k_{\alpha}^{\times}}(\eta_{\alpha})\cdot|V_{\Sigma\alpha}^{\eta^{m_{\alpha}}}|^{\frac{1}{2}}.$
@@
$\Theta_{\omega_{\Sigma\alpha}}([\eta]^{m_{\alpha}})=\operatorname{sgn}_{\mathbb{F}_{p}^{\times}}(-1)^{g_{\alpha}(f_{\alpha}-1)}\cdot\operatorname{sgn}_{k_{\alpha}^{\times}}(\eta_{\alpha})\cdot|V_{\Sigma\alpha}^{\eta^{m_{\alpha}}}|^{\frac{1}{2}}.$
@@ 6.4.2. The case where $\alpha_{\mathrm{res}}$ is symmetric
The action of $\eta]^{m_{\alpha}}$ on $V_{\Sigma\alpha}$ swaps $V_{\Gamma\alpha}$ and $V_{-\Gamma\alpha}$. Since $\eta]^{m_{\alpha}}(X_{\alpha})$ belongs to $V_{\theta^{m_{\alpha}}(\alpha)}=V_{-\sigma_{\alpha}(\alpha)}$, the isomorphism from $V_{\alpha}$ to $V_{-\alpha}$ induced by $\eta]^{m_{\alpha}}$ is $\varsigma_{\alpha}$-linear:
${V_{\alpha}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${V_{\Gamma\alpha}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{[\eta]^{m_{\alpha}}}$${X_{\alpha}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\sum_{\sigma\in\Gamma/\Gamma_{\alpha}}\sigma(X_{\alpha})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${V_{-\alpha}}$${V_{-\Gamma\alpha}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${[\eta]^{m_{\alpha}}\circ\varsigma_{\alpha}(X_{\alpha})}$${\sum_{\sigma\in\Gamma/\Gamma_{\alpha}}[\eta]^{m_{\alpha}}\circ\sigma(X_{\alpha})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
@@
as an element of $\operatorname{Sp}(V_{\Sigma\alpha})\cong\operatorname{Sp}(k_{\alpha}\oplus k_{\alpha})$. Since this automorphism preserves the symplectic form $(x_{+}+x_{-},y_{+}+y_{-})\mapsto\operatorname{Tr}_{k_{\alpha}/\mathbb{F}_{p}}(C\cdot(x_{+}y_{-}-x_{-}y_{+}))$ (see Section 6.1), we must have
@@
To compute $\Theta_{\omega_{\Sigma\alpha}}([\eta]^{m_{\alpha}})$, we utilize Gérardin’s formula (Proposition A.3). Note that $\eta]^{m_{\alpha}}$ is of finite prime-to-$p$ order, hence a semisimple element of $\operatorname{Sp}(V_{\Sigma\alpha})$. We let $T$ be an $\mathbb{F}_{p}$-rational maximal torus of $\operatorname{Sp}(V_{\Sigma\alpha})$ containing $\eta]^{m_{\alpha}}$ (note that such a $T$ may not be unique, so we fix one). Following the notation of Section A.2, we write $P(V_{\Sigma\alpha},T)$ for the set of weights of $T$ in $V_{\Sigma\alpha}$, which is a finite set equipped with an action of $\Sigma_{\mathbb{F}_{p}}:=\pm\Gamma_{\mathbb{F}_{p}}$ (the notion of being symmetric or asymmetric for the weights are defined with respect to this action). Then Proposition A.3 gives
(5) $\Theta_{\omega_{\Sigma\alpha}}([\eta]^{m_{\alpha}})=(-1)^{l(V_{\Sigma\alpha},T;[\eta]^{m_{\alpha}})}\cdot p^{N(V_{\Sigma\alpha};[\eta]^{m_{\alpha}})}\cdot\chi^{T}([\eta]^{m_{\alpha}}).$
@@
  • $l(V_{\Sigma\alpha},T;[\eta]^{m_{\alpha}})\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}|\{\omega\in P(V_{\Sigma\alpha},T)/\Gamma_{\mathbb{F}_{p}}\mid\text{$\epsilon([\eta]^{m_{\alpha}})\neq 1$ for an(y) $\epsilon\in\omega$}\}|$,
  • $N(V_{\Sigma\alpha};[\eta]^{m_{\alpha}})\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\frac{1}{2}\dim_{\mathbb{F}_{p}}{V_{\Sigma\alpha}^{\eta^{m_{\alpha}}}}$ (hence $p^{N(V_{\Sigma\alpha};[\eta]^{m_{\alpha}})}=|V_{\Sigma\alpha}^{\eta^{m_{\alpha}}}|^{\frac{1}{2}}$),
@@
Therefore, our task is to find an $\mathbb{F}_{p}$-rational maximal torus $T$ of $\operatorname{Sp}(V_{\Sigma\alpha})$ containing $\eta]^{m_{\alpha}}$ and compute the set $P(V_{\Sigma\alpha},T)$ including its Galois action. For this, noting that any semisimple conjugacy class of $\operatorname{Sp}(V_{\Sigma\alpha})$ is uniquely determined by the multi-set of eigenvalues (Lemma A.4), we first determine the multi-set of eigenvalues of $\eta]^{m_{\alpha}}$ as an $\mathbb{F}_{p}$-automorphism of $V_{\Sigma\alpha}$.
@@
In the following, we explain an algorithm to pick up an $\mathbb{F}_{p}$-rational maximal torus $T$ of $\operatorname{Sp}(V_{\Sigma\alpha})$ containing $\eta]^{m_{\alpha}}$.
@@
Therefore, if we choose an $\mathbb{F}_{p}$-rational maximal torus $T$ of $\operatorname{Sp}(V_{\Sigma\alpha})$ to be $\prod_{i=1}^{r}k_{i}^{\times}$, then its element $(x_{i})_{i}$ has the same multi-set of eigenvalues as $\eta]^{m_{\alpha}}$ (see Lemma A.5).
The partition of $P(V_{\Sigma\alpha},T)$ into $\Gamma_{\mathbb{F}_{p}}$-orbit is given by
@@
as described in Lemma A.5 (note that each $\Gamma_{\mathbb{F}_{p}}$-orbit is asymmetric). In particular, the number $l(V_{\Sigma\alpha},T;[\eta]^{m_{\alpha}})$ is always even, hence the sign $(-1)^{l(V_{\Sigma\alpha},T;[\eta]^{m_{\alpha}})}$ in the formula (5) is trivial. Since $\epsilon(-)^{\frac{1-q_{\Omega}}{2}}$ for each $\Omega=\{\sigma(-)^{\pm 1}\mid\sigma\in\operatorname{Gal}(k_{i}/\mathbb{F}_{p})\}$ is nothing but the unique quadratic character of $k_{i}^{\times}$, we have $\chi_{T}=\prod_{i=1}^{r}\operatorname{sgn}_{k_{i}^{\times}}$. In summary, we get
$\Theta_{\omega_{\Sigma\alpha}}([\eta]^{m_{\alpha}})=\prod_{i=1}^{r}\operatorname{sgn}_{k_{i}^{\times}}(x_{i})\cdot|V_{\Sigma\alpha}^{\eta^{m_{\alpha}}}|^{\frac{1}{2}}.$
@@
Therefore, if we choose an $\mathbb{F}_{p}$-rational maximal torus $T$ of $\operatorname{Sp}(V_{\Sigma\alpha})$ to be $\prod_{i=1}^{r}k_{i}^{1}\times\prod_{j=1}^{s}(k^{\prime}_{j})^{1}$, then its element $((x_{i})_{i},(x^{\prime}_{j})_{j})$ has the same multi-set of eigenvalues as $\eta]^{m_{\alpha}}$ (see Lemma A.5).
The partition of $P(V_{\Sigma\alpha},T)$ into $\Gamma_{\mathbb{F}_{p}}$-orbit is given by

### v2

@@ 6.3. Descent of the Heisenberg quotient
Recall that $m_{\alpha}$ is the smallest positive integer such that $\theta^{m_{\alpha}}(\alpha)\in\ddot{\alpha}$. We fix an element $\sigma_{\alpha}\in\Gamma_{\pm\alpha_{\mathrm{res}}}/\Gamma_{\alpha}$ as follows:
@@
    When $\alpha$ is asymmetric, we choose $\sigma_{\alpha}$ such that $\theta^{m_{\alpha}}(\alpha)=\pm\sigma_{\alpha}(\alpha)$. Note that, as $\alpha$ is asymmetric, exactly one of $\theta^{m_{\alpha}}(\alpha)$ or $\theta^{m_{\alpha}}(\alpha)$ belongs to $\dot{\alpha}$. In other words, the sign $\pm$” in $\theta^{m_{\alpha}}(\alpha)=\pm\sigma_{\alpha}(\alpha)$ is determined by $\alpha$.
@@
    When $\alpha$ is symmetric, we have $\theta^{m_{\alpha}}(\alpha)\in\ddot{\alpha}=\dot{\alpha}$. Noting this, we choose $\sigma_{\alpha}$ such that $\theta^{m_{\alpha}}(\alpha)=\sigma_{\alpha}(\alpha)$.
@@
Let us show (1); suppose that $\alpha$ is asymmetric. We take any element $\sigma\in\Gamma_{\pm\alpha_{\mathrm{res}}}$ and suppose that $\sigma(\alpha)=\pm\theta_{\mathbf{S}}^{i}(\alpha)$ for some $i$. Then, as $m_{\alpha}$ is the smallest positive integer satisfying $\theta_{\mathbf{S}}^{m_{\alpha}}(\alpha)\in\ddot{\alpha}$, we have $i\equiv jm_{\alpha}\pmod{l_{\alpha}}$ for some $j\in\mathbb{Z}$. Hence we have $\sigma(\alpha)=\sigma_{\alpha}^{j}(\alpha)$. This implies that $\sigma^{-1}\sigma_{\alpha}^{j}$ belongs to $\Gamma_{\alpha}$. In other words, $\sigma_{\alpha}$ generates $\Gamma_{\alpha_{\mathrm{res}}}/\Gamma_{\alpha}$.
@@
$V_{\eta}\cong\bigoplus_{\ddot{\alpha}_{\mathrm{res}}\in\ddot{\Xi}_{\eta}}V_{\eta,\ddot{\alpha}_{\mathrm{res}}},$
@@
$V_{\eta,\ddot{\alpha}_{\mathrm{res}}}\cong\Bigl(\bigoplus_{\ddot{\alpha}^{\prime}\in\Sigma\backslash(\Sigma\times\Theta)\alpha}V_{\ddot{\alpha}^{\prime}}\Bigr)^{\eta}$
for any $\alpha_{\mathrm{res}}\in\Xi_{\eta}$. Let us put $V_{\Theta(\ddot{\alpha})}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\bigoplus_{\ddot{\alpha}^{\prime}\in\Sigma\backslash(\Sigma\times\Theta)\alpha}V_{\ddot{\alpha}^{\prime}}$. In particular, by letting $\Xi_{\mathrm{res}}$ be the set of restricted roots associated to $\Xi$, the set $\Xi_{\eta}$ can be thought of as the set of restricted roots $\alpha_{\mathrm{res}}\in\Xi_{\mathrm{res}}$ such that the $\eta]-$action has a nonzero fixed point in $V_{\Theta(\ddot{\alpha})}$.
@@ 6.4. Twisted characters of Weil representations: asymmetric roots
Let $\alpha\in\Xi_{\mathrm{asym}}$. We compute $\Theta_{\omega_{\ddot{\alpha}}}([\eta]^{m_{\alpha}})$, which constitutes the right-hand side of Corollary 6.6. Recall that $V_{\ddot{\alpha}}=V_{\dot{\alpha}}\oplus V_{-\dot{\alpha}}\cong V_{\alpha}\oplus V_{-\alpha}$, where $V_{\alpha}$ and $V_{-\alpha}$ are $1$-dimensional $k_{\alpha}$-vector spaces, which are identified with $k_{\alpha}$ by fixing nonzero elements $X_{\alpha}\in V_{\alpha}$ and $X_{-\alpha}\in V_{-\alpha}$ (see Section 6.1).
We carry out a case-by-case computation on the symmetry of $\alpha_{\mathrm{res}}$. Recall that we have exactly one of $\theta^{m_{\alpha}}(\alpha)\in\dot{\alpha}$ or $\theta^{m_{\alpha}}(\alpha)\in-\dot{\alpha}$; we have fixed a $\sigma_{\alpha}$ satisfying $\theta^{m_{\alpha}}(\alpha)=\sigma_{\alpha}(\alpha)$ or $\sigma_{\alpha}(\alpha)$ according to it. In the former case where $\theta^{m_{\alpha}}(\alpha)=\sigma_{\alpha}(\alpha)$, $\sigma_{\alpha}$ belongs to $\Gamma_{\alpha_{\mathrm{res}}}$, which implies that $\alpha_{\mathrm{res}}$ is asymmetric. On the other hand, in the latter case where $\theta^{m_{\alpha}}(\alpha)=-\sigma_{\alpha}(\alpha)$, the Galois group of $F_{\alpha}/F_{\alpha_{\mathrm{res}}}$ is a proper subgroup of that of $F_{\alpha}/F_{\pm\alpha_{\mathrm{res}}}$ generated by $\sigma_{\alpha}^{2}$, which implies that $\alpha_{\mathrm{res}}$ is symmetric. Thus, also recalling Lemma 6.8, we have the following possibilities:
@@ 6.4.1. The case where $\alpha_{\mathrm{res}}$ is asymmetric
The action of $\eta]^{m_{\alpha}}$ on $V_{\ddot{\alpha}}$ preserves $V_{\dot{\alpha}}$ and $V_{-\dot{\alpha}}$. Since $\eta]^{m_{\alpha}}(X_{\alpha})$ belongs to $V_{\theta^{m_{\alpha}}(\alpha)}=V_{\sigma_{\alpha}(\alpha)}$, the induced action of $\eta]^{m_{\alpha}}$ on $V_{\alpha}$ is $\varsigma_{\alpha}$-linear:
${V_{\alpha}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${V_{\dot{\alpha}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{[\eta]^{m_{\alpha}}}$${X_{\alpha}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\sum_{\sigma\in\Gamma/\Gamma_{\alpha}}\sigma(X_{\alpha})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${V_{\alpha}}$${V_{\dot{\alpha}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${[\eta]^{m_{\alpha}}\circ\varsigma_{\alpha}(X_{\alpha})}$${\sum_{\sigma\in\Gamma/\Gamma_{\alpha}}[\eta]^{m_{\alpha}}\circ\sigma(X_{\alpha})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
@@
as an element of $\operatorname{Sp}(V_{\ddot{\alpha}})\cong\operatorname{Sp}(k_{\alpha}\oplus k_{\alpha})$. Hence, by Corollary A.8, we get
$\Theta_{\omega_{\ddot{\alpha}}}([\eta]^{m_{\alpha}})$ $=\operatorname{sgn}_{\mathbb{F}_{p}^{\times}}(\det(\eta_{\alpha}\circ\varsigma_{\alpha}\mid k_{\alpha}))\cdot|V_{\ddot{\alpha}}^{\eta^{m_{\alpha}}}|^{\frac{1}{2}}$
$=\operatorname{sgn}_{\mathbb{F}_{p}^{\times}}(\det(\varsigma_{\alpha}\mid k_{\alpha}))\cdot\operatorname{sgn}_{k_{\alpha}^{\times}}(\eta_{\alpha})\cdot|V_{\ddot{\alpha}}^{\eta^{m_{\alpha}}}|^{\frac{1}{2}}.$
@@
$\Theta_{\omega_{\ddot{\alpha}}}([\eta]^{m_{\alpha}})=\operatorname{sgn}_{\mathbb{F}_{p}^{\times}}(-1)^{g_{\alpha}(f_{\alpha}-1)}\cdot\operatorname{sgn}_{k_{\alpha}^{\times}}(\eta_{\alpha})\cdot|V_{\ddot{\alpha}}^{\eta^{m_{\alpha}}}|^{\frac{1}{2}}.$
@@ 6.4.2. The case where $\alpha_{\mathrm{res}}$ is symmetric
The action of $\eta]^{m_{\alpha}}$ on $V_{\ddot{\alpha}}$ swaps $V_{\dot{\alpha}}$ and $V_{-\dot{\alpha}}$. Since $\eta]^{m_{\alpha}}(X_{\alpha})$ belongs to $V_{\theta^{m_{\alpha}}(\alpha)}=V_{-\sigma_{\alpha}(\alpha)}$, the isomorphism from $V_{\alpha}$ to $V_{-\alpha}$ induced by $\eta]^{m_{\alpha}}$ is $\varsigma_{\alpha}$-linear:
${V_{\alpha}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${V_{\dot{\alpha}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{[\eta]^{m_{\alpha}}}$${X_{\alpha}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\sum_{\sigma\in\Gamma/\Gamma_{\alpha}}\sigma(X_{\alpha})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${V_{-\alpha}}$${V_{-\dot{\alpha}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${[\eta]^{m_{\alpha}}\circ\varsigma_{\alpha}(X_{\alpha})}$${\sum_{\sigma\in\Gamma/\Gamma_{\alpha}}[\eta]^{m_{\alpha}}\circ\sigma(X_{\alpha})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
@@
as an element of $\operatorname{Sp}(V_{\ddot{\alpha}})\cong\operatorname{Sp}(k_{\alpha}\oplus k_{\alpha})$. Since this automorphism preserves the symplectic form $(x_{+}+x_{-},y_{+}+y_{-})\mapsto\operatorname{Tr}_{k_{\alpha}/\mathbb{F}_{p}}(C\cdot(x_{+}y_{-}-x_{-}y_{+}))$ (see Section 6.1), we must have
@@
To compute $\Theta_{\omega_{\ddot{\alpha}}}([\eta]^{m_{\alpha}})$, we utilize Gérardin’s formula (Proposition A.3). Note that $\eta]^{m_{\alpha}}$ is of finite prime-to-$p$ order, hence a semisimple element of $\operatorname{Sp}(V_{\ddot{\alpha}})$. We let $T$ be an $\mathbb{F}_{p}$-rational maximal torus of $\operatorname{Sp}(V_{\ddot{\alpha}})$ containing $\eta]^{m_{\alpha}}$ (note that such a $T$ may not be unique, so we fix one). Following the notation of Section A.2, we write $P(V_{\ddot{\alpha}},T)$ for the set of weights of $T$ in $V_{\ddot{\alpha}}$, which is a finite set equipped with an action of $\Sigma_{\mathbb{F}_{p}}:=\pm\Gamma_{\mathbb{F}_{p}}$ (the notion of being symmetric or asymmetric for the weights are defined with respect to this action). Then Proposition A.3 gives
(5) $\Theta_{\omega_{\ddot{\alpha}}}([\eta]^{m_{\alpha}})=(-1)^{l(V_{\ddot{\alpha}},T;[\eta]^{m_{\alpha}})}\cdot p^{N(V_{\ddot{\alpha}};[\eta]^{m_{\alpha}})}\cdot\chi^{T}([\eta]^{m_{\alpha}}).$
@@
  • $l(V_{\ddot{\alpha}},T;[\eta]^{m_{\alpha}})\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}|\{\omega\in P(V_{\ddot{\alpha}},T)/\Gamma_{\mathbb{F}_{p}}\mid\text{$\epsilon([\eta]^{m_{\alpha}})\neq 1$ for an(y) $\epsilon\in\omega$}\}|$,
  • $N(V_{\ddot{\alpha}};[\eta]^{m_{\alpha}})\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\frac{1}{2}\dim_{\mathbb{F}_{p}}{V_{\ddot{\alpha}}^{\eta^{m_{\alpha}}}}$ (hence $p^{N(V_{\ddot{\alpha}};[\eta]^{m_{\alpha}})}=|V_{\ddot{\alpha}}^{\eta^{m_{\alpha}}}|^{\frac{1}{2}}$),
@@
Therefore, our task is to find an $\mathbb{F}_{p}$-rational maximal torus $T$ of $\operatorname{Sp}(V_{\ddot{\alpha}})$ containing $\eta]^{m_{\alpha}}$ and compute the set $P(V_{\ddot{\alpha}},T)$ including its Galois action. For this, noting that any semisimple conjugacy class of $\operatorname{Sp}(V_{\ddot{\alpha}})$ is uniquely determined by the multi-set of eigenvalues (Lemma A.4), we first determine the multi-set of eigenvalues of $\eta]^{m_{\alpha}}$ as an $\mathbb{F}_{p}$-automorphism of $V_{\ddot{\alpha}}$.
@@
In the following, we explain an algorithm to pick up an $\mathbb{F}_{p}$-rational maximal torus $T$ of $\operatorname{Sp}(V_{\ddot{\alpha}})$ containing $\eta]^{m_{\alpha}}$.
@@
Therefore, if we choose an $\mathbb{F}_{p}$-rational maximal torus $T$ of $\operatorname{Sp}(V_{\ddot{\alpha}})$ to be $\prod_{i=1}^{r}k_{i}^{\times}$, then its element $(x_{i})_{i}$ has the same multi-set of eigenvalues as $\eta]^{m_{\alpha}}$ (see Lemma A.5).
The partition of $P(V_{\ddot{\alpha}},T)$ into $\Gamma_{\mathbb{F}_{p}}$-orbit is given by
@@
as described in Lemma A.5 (note that each $\Gamma_{\mathbb{F}_{p}}$-orbit is asymmetric). In particular, the number $l(V_{\ddot{\alpha}},T;[\eta]^{m_{\alpha}})$ is always even, hence the sign $(-1)^{l(V_{\ddot{\alpha}},T;[\eta]^{m_{\alpha}})}$ in the formula (5) is trivial. Since $\epsilon(-)^{\frac{1-q_{\Omega}}{2}}$ for each $\Omega=\{\sigma(-)^{\pm 1}\mid\sigma\in\operatorname{Gal}(k_{i}/\mathbb{F}_{p})\}$ is nothing but the unique quadratic character of $k_{i}^{\times}$, we have $\chi_{T}=\prod_{i=1}^{r}\operatorname{sgn}_{k_{i}^{\times}}$. In summary, we get
$\Theta_{\omega_{\ddot{\alpha}}}([\eta]^{m_{\alpha}})=\prod_{i=1}^{r}\operatorname{sgn}_{k_{i}^{\times}}(x_{i})\cdot|V_{\ddot{\alpha}}^{\eta^{m_{\alpha}}}|^{\frac{1}{2}}.$
@@
Therefore, if we choose an $\mathbb{F}_{p}$-rational maximal torus $T$ of $\operatorname{Sp}(V_{\ddot{\alpha}})$ to be $\prod_{i=1}^{r}k_{i}^{1}\times\prod_{j=1}^{s}(k^{\prime}_{j})^{1}$, then its element $((x_{i})_{i},(x^{\prime}_{j})_{j})$ has the same multi-set of eigenvalues as $\eta]^{m_{\alpha}}$ (see Lemma A.5).
The partition of $P(V_{\ddot{\alpha}},T)$ into $\Gamma_{\mathbb{F}_{p}}$-orbit is given by